A stationary store has decided to accept a large shipment of ball-point pens if an inspection of 15 randomly selected pens yields no more than two defective pens.

(a) Find the probability that this shipment is accepted if 5% of the total shipment is defective. (Use 3 decimal places.)

(b) Find the probability that this shipment is not accepted if 15% of the total shipment is defective. (Use 3 decimal places.)
Answer #1

n = 15

P(defective) = 0.05

a) We have to find:

P(x ≤ 2) = P(x = 2) + P(x = 1) + P(x = 0)

P(x ≤ 2) = 0.964

b) P(defective) = 0.15

n = 15

We need to find:

P(x >2) = 0.396